# Diverse Score Distillation

### v1

@@ 1 Introduction
In this work, we present an alternate formulation for distilling diffusion models that overcomes this limitation, and allows diverse (3D) generation. Our approach is inspired by the ODE perspective on sampling from diffusion models [30] which highlights that a trained model can be viewed as inducing a learned ODE that maps noise samples to data, and different starting points for the ODEs (i.e. different initial noise samples) yield diverse data samples. Building on this insight, we derive a gradient that allows an optimization to ‘follow’ an ODE, and enables diverse optimization outputs by simply specifying differing ODE (noise) initializations. We first validate our approach for 2D image generation via optimization and show that, akin to DDIM sampling, it allows diverse and high-fidelity generation.
@@ 3 Method
Our goal is to develop a diffusion-guided optimization framework that ensures diversity in the optimized outputs i.e. multiple instantiations of the optimization process (using different seeds) should result in different output samples, thus allowing, for example, multiple plausible 3D representations given text conditioning. To develop such a formulation, we take inspiration from the sampling processing denoising diffusion models which inherently yields such diverse samples. We present an initial formulation (Sec. 3.2) that seeks to enforce the optimization to track the evolution of a diffusion sampling process. In Sec. 3.3, we then extend this to incorporate scenarios where this might not be exactly feasible (e.g. the renderings of a 3D representation being optimized cannot perfectly match target 2D diffusion processes). Before detailing our formulation, we first review some preliminaries about denoising diffusion, score distillation, and DDIM sampling, all of which play a central role in our approach.
@@ 3.1 Background
This update actually corresponds to a discretized approximation for the ‘Probability Flow Ordinary Differential Equation’ (PF-ODE)  [30] for the diffusion model:
@@
Data and Noise Trajectories Induced by DDIM.
While the DDIM process follows (ODE-based) trajectories for $\mathbf{x}(t)$, a recent score distillation approach [19] noted that these also induce trajectories in denoised data and predicted noise space. In particular, given $\mathbf{x}(t)$, we can define a trajectory for the corresponding predicted noise $\epsilon(t)=\epsilon_{\theta}^{t}(\mathbf{x}(t),\mathbf{y})$ and a ‘one-step data prediction’ $\mathbf{x}_{0}(t)$ that satisfies:
$\mathbf{x}(t)=\sqrt{\alpha(t)}~{}\mathbf{x}_{0}(t)+\sqrt{1-\alpha(t)}~{}\epsilon(t)$ (3)
One can also derive a discrete update for the one-step data prediction trajectory:
@@
We visualize the ODE trajectory for $\mathbf{x}(t)$ and the induced ones for $\epsilon(t)$ and $\mathbf{x}_{0}(t)$ in Fig. 2.
@@ 3.2 Sampling-based Score Distillation
To leverage pre-trained diffusion models for guiding optimization e.g. of a 3D representation with parameters $\psi$ given a text prompt $\mathbf{y}$, a score distillation based approach relies on obtaining gradients for an image e.g. a rendering $\mathbf{x}_{\pi}\equiv g(\psi,\pi)$. Specifically, a diffusion-based score functions outputs a gradient $\nabla\mathbf{x}_{\pi}$ (by adding some noise corresponding to a random/scheduled timestep $t$), and this process can used to optimize the 3D representation:
The central design choice in this process is how one can obtain the gradient $\nabla\mathbf{x}_{\pi}$ (as well as how one can sample the diffusion timestep $t$ for computing the score).
To motivate our formulation, we first consider a simpler optimization scenario where we directly optimizing a 2D representation $\mathbf{x}_{\pi}$ using gradients $\nabla\mathbf{x}_{\pi}$. Our insight is that we can define a gradient update that allows the optimized variable $\mathbf{x}_{\pi}$, in the ideal case, to follow an ODE-based trajectory. Specifically, given a random seed $\epsilon^{*}$, we can ensure that the $\mathbf{x}_{\pi}$ evolves akin to the ‘one-step data prediction’ $\mathbf{x}_{0}(t)$ for the ODE trajectory specified by $\mathbf{x}(T)=\epsilon^{*}$. Such an update would ensure generation fidelity as the final image would be likely under the diffusion model (as it corresponds to on ODE output), while also allowing diversity by simply sampling different seeds $\epsilon^{*}\sim\mathcal{N}(0,\mathbf{I})$. To operationalize this, we follow an annealed time scheduled (varying $t$ during optimization steps from $T$ to 00), and define the gradient for the variable $\mathbf{x}_{\pi}$ to mimic the update defined earlier (Eq. 4):
@@
$\nabla_{\mathbf{x}_{\pi}}^{Sampling}:=\sigma(t-\delta)[\epsilon_{\theta}^{t-\delta}(\mathbf{x}(t-\delta),\mathbf{y})-\epsilon_{\theta}^{t}(\mathbf{x}(t),\mathbf{y})]$
Figure 4: 2D Distillation Results using DDIM as reference. The prompts are “A hamburger”, “A dragon with flames coming out of its mouth” and “An apple”. We assign a fixed initial noise for each grid. When the number of DDIM steps equals the optimization step, our method (DMD*) resembles the DDIM sample. When different, the optimized image will deviate from the original ODE by a small margin. We observe that DSD yields more diverse and plausible generations compared to alternates.
We refer to the formulation in Sec. 3.2 as ‘Sampling-based Score Distillation’ as the resulting optimization simulates the ODE sampling process from $\mathbf{x}(T)=\epsilon^{*}$. In fact, when the number of optimization steps are chosen to be exactly equal to DDIM sampling steps, outputs from this formulation are equivalent to DDIM samples, and we discuss in the appendix how minor modifications to the step size and $\delta$ can yield similar generations if they are not.
@@ 3.3 Distillation via Interpolation Approximation
While we motivated the formulation in Sec. 3.2 via optimization of a 2D representation, we can use the derived score function to also optimize a 3D representation. To this end, we can choose a common seed $\epsilon^{*}$ across all views of the 3D representation (see appendix for empirical justification) and a separate view-conditioned prompt $\mathbf{y}_{\pi}$ [1] to instantiate the ODE that each view $\mathbf{x}_{\pi}$ should follow. To obtain different 3D shapes $\psi_{i}$ and $\psi_{j}$, we can choose $\epsilon^{*}_{i}\neq\epsilon^{*}_{j}$.
@@
To address this issue, we observe that both $\mathbf{x}(t)$ and $\mathbf{x}(t-\delta)$ in Sec. 3.2 can be expressed as linear combinations of $\mathbf{x}_{0}(t)$ and $\epsilon(t)$. Our key insight is that we can instead re-define them to be an interpolation of $\mathbf{x}_{\pi}$ and $\epsilon(t)$. In the ideal case, $\mathbf{x}_{\pi}$ evolves similarly to $\mathbf{x}_{0}(t)$, but when it does not, this alteration allows a correction mechanism as the gradients are dependent on the current optimization variable.
We thus generalize our previous formulation for scenarios where the optimization may drift from the ODE (shown in Fig. 3 ). This leads to our Diverse Score Distillation (DSD) gradient $\nabla_{\mathbf{x}_{\pi}}^{DSD}=\texttt{{DSD}}(\epsilon^{*},\mathbf{y},t,\delta,\mathbf{x}_{\pi})$ as:
@@
$\hat{\mathbf{x}}(t)=\sqrt{\alpha(t)}\mathbf{x}_{\pi}(t)+\sqrt{1-\alpha(t)}\epsilon(t)$ (7)
$\hat{\mathbf{x}}(t-\delta)=\sqrt{\alpha(t-\delta)}\mathbf{x}_{\pi}(t)+\sqrt{1-\alpha(t)}\epsilon(t)$
$\nabla_{\mathbf{x}_{\pi}}^{DSD}=\sigma(t-\delta)[\epsilon_{\theta}^{t-\delta}(\hat{\mathbf{x}}(t-\delta),\mathbf{y})-\epsilon_{\theta}^{t}(\hat{\mathbf{x}}(t),\mathbf{y})].$
Using this formulation, we can instantiate our DSD-based 3D optimization procedure in Algorithm 1.
@@
6:     $\mathbf{x}_{\pi}(t)\leftarrow g(\psi,\pi)$
7:     $\nabla_{\mathbf{x}_{\pi}}\leftarrow\texttt{{DSD}}(\epsilon^{*},\mathbf{y},t,\delta,\mathbf{x}_{\pi}(t))$
8:     $\nabla_{\psi}=w(t)\nabla_{\mathbf{x}_{\pi}}\frac{\partial g}{\partial\psi}$
@@
SDS  [23] (Random $t$ ) ASD [20] (Random $t$)
$\epsilon\sim\mathcal{N}(0,\mathbf{I})$ $\epsilon\sim\mathcal{N}(0,\mathbf{I})$
$\mathbf{x}(t-\delta)=\sqrt{\alpha(t-\delta)}\mathbf{x}_{\pi}+\sqrt{1-\alpha(t-\delta)}\epsilon$ $\mathbf{x}(t)=\sqrt{\alpha(t)}\mathbf{x}_{\pi}+\sqrt{1-\alpha(t)}\epsilon$
$\nabla_{\mathbf{x}_{\pi}}=w(t-\delta)[\epsilon_{\theta}^{t-\delta}(\mathbf{x}(t-\delta),\mathbf{y})-\epsilon]$ $\mathbf{x}(t-\delta)=\sqrt{\alpha(t-\delta)}\mathbf{x}_{\pi}+\sqrt{1-\alpha(t-\delta)}\epsilon$
$\nabla_{\mathbf{x}_{\pi}}=w(t-\delta)[\epsilon_{\theta}^{t-\delta}(\mathbf{x}(t-\delta),\mathbf{y})-\epsilon_{\theta}^{t}(\mathbf{x}(t),\mathbf{y})]$
SDI  [19] (Annealed $t$) Consistent 3D  [34] (Annealed $t$)
$\mathbf{x}(t)=\texttt{DDIM\_Inversion}(\mathbf{x}_{\pi},0\rightarrow t,\mathbf{y})$ $\epsilon^{*}=\texttt{Fixed}$
$\epsilon(t)=\epsilon_{\theta}^{t}(\mathbf{x}(t),\mathbf{y})$ $\mathbf{x}(t)=\sqrt{\alpha(t)}\mathbf{x}_{\pi}+\sqrt{1-\alpha(t)}\epsilon$
$\mathbf{x}(t-\delta)=\sqrt{\alpha(t-\delta)}\mathbf{x}_{\pi}+\sqrt{1-\alpha(t-\delta)}\epsilon(t)$ $\mathbf{x}_{0}(t)=\left(\mathbf{x}(t)-\sqrt{1-\alpha(t)}\epsilon_{\theta}^{t}(\mathbf{x}(t),\mathbf{y})\right)/\sqrt{\alpha(t)}$
$\nabla_{\mathbf{x}_{\pi}}=w(t-\delta)[\epsilon_{\theta}^{t-\delta}(\mathbf{x}(t-\delta),\mathbf{y})-\epsilon_{\theta}^{t}(\mathbf{x}(t),\mathbf{y})]$ $\mathbf{x}(t-\delta)=\sqrt{\alpha(t-\delta)}\mathbf{x}_{0}(t)+\sqrt{1-\alpha(t-\delta)}\epsilon_{\theta}^{t}(\mathbf{x}(t),\mathbf{y})$
$\nabla_{\mathbf{x}_{\pi}}={\color[rgb]{1,0,0}-}w(t-\delta)[\epsilon_{\theta}^{t-\delta}(\mathbf{x}(t-\delta),\mathbf{y})-\epsilon_{\theta}^{t}(\mathbf{x}(t),\mathbf{y})]$
Table 1: A unifying perspective on score distillation. The mentioned methods can be thought of as instantiations of Eq. 8. This allows us to analyze these methods from the ODE perspective. Note that besides loss formulation, the schedule for diffusion noise level $t$ can be varying across methods.
4 An ODE Perspective on Score Distillation
Our formulation in Sec. 3.3 computes a gradient of the form:
$\nabla_{\mathbf{x}_{\pi}}=w(t-\delta)[\epsilon_{\theta}^{t-\delta}(\mathbf{x}(t-\delta)\mathbf{y})-\epsilon_{\theta}^{t}(\mathbf{x}(t),\mathbf{y})].$ (8)
We notice that gradient computation in several prior score distillation approaches can be similarly formulated as the difference between two (predicted) noises, with key differences in the variables input for this prediction. As our work derives these updates from an ODE sampling perspective, we examine the prior formulations under this lens to explain their properties.
Score Distillation Sampling (SDS) [23] samples a random $\epsilon$ at each iteration, implying that it would get high-variance gradients, thus preventing diverse generations (which require consistent low-variance gradients). Moreover, the term $\epsilon$ in the gradient can be considered as ‘noise prediction’ at $t=T$, and is thus a ‘coarser’ gradient than the one from noise prediction using $\mathbf{x}(t)$, which accounts for the overly smooth results observed in SDS.
@@ 4 An ODE Perspective on Score Distillation
Consistent 3D [34] attempts to use a fixed ODE throughout each optimization process by setting a constant $\epsilon^{*}$. However, the linear interpolation between $\mathbf{x}_{\pi}$ and $\epsilon^{*}$ is an approximation of the current ODE (unlike our approach which uses DDIM to solve the ODE). More surprisingly however, the gradient applied by Consistent 3D is not similar to Eq. 8; it is actually the opposite! In practice, their approach requires a large $\delta$ (and classifier-free guidance) to synthesis plausible generations, with the intuition that noise predictions from larger timesteps (under high CFG) can point to likely images given text conditioning. While the different seeds $\epsilon^{*}$ can yield diverse generations, the fidelity of their results suffers perhaps due to the inconsistency of the gradient update.
@@ 5 Experiments
We seek to empirically validate the ability of DSD to guide diffusion-based optimization and generated diverse and high-fidelity outputs. We first analyze our formulation in the simple but informative task of guiding text-based image optimization and then examine text-to-3D generation. We further ablate some key decision decisions and also highlight the applicability for distillation-based single-view reconstruction.
Figure 5: Generation Comparison. Text-to-3D generations, with each method generating 4 samples. We find that DSD is capable of generating higher-quality 3D shapes while being much more diverse. It is also robust to harder prompts.
@@
We select SDS  [23], HiFA [38] and ASD  [20] to represent the methods that sample random noise, and select consistent 3D  [34] and SDI  [19] as the approaches that perform score distillation from an ODE perspective. We also include VSD  [33] as it minimizes the KL divergence between the rendering and image distribution, thus promising diverse generation. For fair comparison across methods, we adopt the codebase from SDI (which builds on a commonly adopted implementation [6]) and only vary the score function across methods (except for VSD, where we use its released code for multi-particle optimization).

### v2

@@ 1 Introduction
In this work, we present an alternate formulation for distilling diffusion models that overcomes this limitation, and allows diverse (3D) generation. Our approach is inspired by the ODE perspective on sampling from diffusion models [31] which highlights that a trained model can be viewed as inducing a learned ODE that maps noise samples to data, and different starting points for the ODEs (i.e. different initial noise samples) yield diverse data samples. Building on this insight, we derive a gradient that allows an optimization to ‘follow’ an ODE, and enables diverse optimization outputs by simply specifying different ODE (noise) initializations. We first validate our approach for 2D image generation via optimization and show that, akin to DDIM sampling, it allows diverse and high-fidelity generation.
@@ 3 Method
Our goal is to develop a diffusion-guided optimization framework that ensures diversity in the optimized outputs i.e. multiple instantiations of the optimization process (using different seeds) should result in different output samples, thus allowing, for example, multiple plausible 3D representations given text conditioning. To develop such a formulation, we take inspiration from the sampling process in denoising diffusion models which inherently yields such diverse samples. We present an initial formulation (Sec. 3.2) that seeks to enforce the optimization to track the evolution of a diffusion sampling process. In Sec. 3.3, we then extend this to incorporate scenarios where this might not be exactly feasible (e.g. the renderings of a 3D representation being optimized cannot perfectly match target 2D diffusion processes). Before decribing our formulation, we first review some preliminaries about denoising diffusion, score distillation, and DDIM sampling, all of which play a central role in our approach. After outlining our approach, we place our formulation in context of prior score distillation methods, in particular focusing on the diversity of sampling and relation to ODEs (Sec. 3.4).
@@ 3.1 Background
, where $\mathbf{y}$ is a conditioning variable (CLIP text embeddings for the case of text-to-image generation). This update actually corresponds to a discretized approximation for the ‘Probability Flow Ordinary Differential Equation’ (PF-ODE)  [31] for the diffusion model:
@@
Figure 4: 2D Distillation Results using DDIM as reference. The prompts are “A hamburger”, “A dragon with flames coming out of its mouth” and “An apple”. We assign a fixed initial noise for each grid. When the number of DDIM steps equals the optimization step, our method (DSD*) resembles the DDIM sample. When different, the optimized image will deviate from the original ODE by a small margin. We observe that DSD yields more diverse and plausible generations compared to alternates.
Data and Noise Trajectories Induced by DDIM. We can define a ‘one-step data prediction’ variable as:
$\mathbf{x}_{0}(t)=\bar{\mathbf{x}}(t)-\sigma(t)\epsilon(t)$ (3)
, where $\epsilon(t)=\epsilon_{\theta}^{t}(\mathbf{x}(t),\mathbf{y})$. Combining Eq. 2 and Eq. 3, we can define an induced ODE in $\mathbf{x}_{0}(t)$ space:
$\frac{d\mathbf{x}_{0}(t)}{dt}=-\sigma(t)\frac{d}{dt}\epsilon_{\theta}^{t}(\mathbf{x}(t),\mathbf{y}).$ (4)
A discrete update for this one-step data prediction trajectory can also be derived as:
@@
As noticed in  [20, 15], the updates in $\mathbf{x}_{0}$ space have smaller variance compared to that of $\mathbf{x}_{t}$, which is desirable for an optimization process. We visualize the ODE trajectory for $\mathbf{x}(t)$ and the induced ones for $\epsilon(t)$ and $\mathbf{x}_{0}(t)$ in Fig. 2.
Score Distillation. To leverage pre-trained diffusion models for guiding optimization, a score distillation-based approach relies on obtaining gradients for an image. We use $\mathbf{x}_{\pi}$ to represent the image whose gradient ($\nabla\mathbf{x}_{\pi}$) will be used to update the optimizable parameter $\psi$. For our 2D experiments, where we directly optimize an image, $x_{\pi}=\psi$ and $\partial{g}/\partial{\psi}=I$. For 3D generation, $\psi$ denotes the parameter of the 3D shape, and $x_{\pi}$ is the rendering of 2D images over camera poses $\pi$, i.e. $x_{\pi}=g(\psi,\pi)$. Here, $g$ is the differential renderer. Using the the chain rule, one can get:
, where $w(t)$ is a weighting term to normalize gradients at various noise levels for stable optimization. The central design choice in this process is how one can obtain the gradient $\nabla\mathbf{x}_{\pi}$ (as well as how to sample the diffusion timestep $t$ for computing the score).
@@ 3.2 Sampling-based Score Distillation
For simplicity, we first consider a 2D image generation scenario i.e. using score distillation to optimize an image given a text description ($\psi\equiv\mathbf{x}_{\pi}$). We seek to define a gradient update that allows the optimized variable $\psi=\mathbf{x}_{\pi}$ to follow a $\mathbf{x}_{0}$ ODE trajectory (Eq. 4). Such an update would ensure generation fidelity, as the final image would be likely under the diffusion model (as it corresponds to on ODE output), while also allowing diversity by simply sampling different seeds $\epsilon^{*}\sim\mathcal{N}(0,\mathbf{I})$ as starting points for the ODE. We can define such a gradient by adapting the update rule for $\mathbf{x}_{0}(t)$ (Eq. 5):
$\nabla_{\mathbf{x}_{\pi}}^{\text{Sampling}}:=\sigma(t-\delta)[\epsilon_{\theta}^{t-\delta}(\mathbf{x}(t-\delta),\mathbf{y})-\epsilon_{\theta}^{t}(\mathbf{x}(t),\mathbf{y})]$ (7)
To ensure that $\nabla_{\mathbf{x}_{\pi}}^{\text{Sampling}}$ allows the update of $\mathbf{x}_{\pi}$ to follow one underlying ODE (specified by a random seed $\epsilon^{*}$), both $\mathbf{x}(t-\delta)$ and $\mathbf{x}(t)$ should satisfy the DDIM PF-ODE. As shown in the appendix, we guarantee this by defining:
@@
[…]
We refer to the formulation in Eq. 7 and Sec. 3.2 as ‘Sampling-based Score Distillation’ as the resulting optimization simulates the ODE sampling process from $\mathbf{x}(T)=\epsilon^{*}$. In fact, when the number of optimization steps are chosen to be exactly equal to DDIM sampling steps, outputs from this formulation are equivalent to DDIM samples, and we discuss in the appendix how minor modifications to the step size and $\delta$ can yield similar generations if they are not.
@@ 3.3 Distillation via Interpolation Approximation
While we motivated the formulation in Sec. 3.2 via optimization of a 2D representation, we can use the derived score function to also optimize a 3D representation. As shown by Perp-Neg [1], images generated from the same seeds with view-conditioned prompts using the orthogonal update rule can be treated as the set of images from one object. Therefore, we use a common seed $\epsilon^{*}$ across all views of the 3D representation (see appendix for empirical justification) with separate view-conditioned prompt $\mathbf{y}_{\pi}$ [1] to instantiate the ODE that each view $\mathbf{x}_{\pi}$ should follow. To obtain different 3D shapes $\psi_{i}$ and $\psi_{j}$, we can choose $\epsilon^{*}_{i}\neq\epsilon^{*}_{j}$.
@@
To address this issue, we observe that both $\mathbf{x}(t)$ and $\mathbf{x}(t-\delta)$ in Sec. 3.2 can be expressed as linear combinations of $\mathbf{x}_{0}(t)$ and $\epsilon(t)$. Our key insight is that we can instead re-define them to be an interpolation of $\mathbf{x}_{\pi}$ and $\epsilon(t)$:
@@
$\hat{\mathbf{x}}(t)=\sqrt{\alpha(t)}\mathbf{x}_{\pi}+\sqrt{1-\alpha(t)}\epsilon(t)$ (9)
$\hat{\mathbf{x}}(t-\delta)=\sqrt{\alpha(t-\delta)}\mathbf{x}_{\pi}+\sqrt{1-\alpha(t)}\epsilon(t)$
In the ideal case, $\mathbf{x}_{\pi}$ evolves similarly to $\mathbf{x}_{0}(t)$, but when it does not, this alteration allows a correction mechanism as the gradients are dependent on the current optimization variable. We thus generalize our previous formulation for scenarios where the optimization may drift from the ODE (shown in Fig. 3 ). This leads to our Diverse Score Distillation (DSD) gradient:
$\nabla_{\mathbf{x}_{\pi}}^{DSD}=\sigma(t-\delta)[\epsilon_{\theta}^{t-\delta}(\hat{\mathbf{x}}(t-\delta),\mathbf{y})-\epsilon_{\theta}^{t}(\hat{\mathbf{x}}(t),\mathbf{y})]$ (10)
, where we obtain $\hat{\mathbf{x}}(t-\delta)$ and $\hat{\mathbf{x}}(t)$ from Sec. 3.3. Using this formulation, we can instantiate our DSD-based 3D optimization procedure in Algorithm 1.
@@
1:  Initialization: 3D parameter $\psi$, trained text-to-image diffusion model $\epsilon_{\theta}^{t}$, prompt y, set of cameras around 3D shape $\Pi$, differentiable renderer $g$
@@
6:     $\mathbf{x}_{\pi}\leftarrow g(\psi,\pi)$
7:     Compute $\nabla_{\mathbf{x}_{\pi}}^{\text{DSD}}$ (Eq. 10)
8:     $\nabla_{\psi}=w(t)\nabla_{\mathbf{x}_{\pi}}^{\text{DSD}}\frac{\partial g}{\partial\psi}$
@@
3.4 Comparing Score Distillation Formulations
Our approach operationalized two key insights: a) if the optimization approach follows the evolution of an ODE, we can expect higher fidelity outputs as these are likely under the diffusion model, and b) the ability to control the generation process (via a random seed) can yield diverse generations. Below, we briefly review alternate score distillation frameworks, in particular highlighting whether they follow an ODE and/or can generate diverse output, and also summarize this in in Tab. 1.
Score Distillation Sampling (SDS).
SDS defines $\nabla_{\mathbf{x}_{\pi}}^{\text{SDS}}=\mathbb{E}_{\epsilon}[\epsilon_{\theta}^{t}(\mathbf{x}(t),\mathbf{y})-\epsilon]$, where $\mathbf{x}(t)$ is defined by adding random noise to $\mathbf{x}_{\pi}$: $\mathbf{x}(t)=\sqrt{\alpha(t)}\mathbf{x}_{\pi}+\sqrt{1-\alpha(t)}\epsilon$. While this corresponds to the gradient of image log-likelihood under diffusion, the gradient update does not correspond to an ODE. Moreover, the randomly sampled $t$ and $\epsilon$ make this a high-variance estimate.
@@ 3.4 Comparing Score Distillation Formulations
Consistent3D [36] attempts to follow the consistency ODE [32] by setting a constant $\epsilon^{*}$. However, this does not define a valid ODE in the space of denoising diffusion models (including StableDiffusion), necessitating a large classifier-free guidance (CFG) [9] scale (50-100) for the method to be effective, which in turn reduces diversity. In contrast, our approach operates with a CFG of 7.5, aligning with the standard 2D sampling process.
SDS ASD SDI Consistent 3D VSD DSD
Diversity
ODE-following
Table 1: Conceptual Comparisons of Score Distillation Methods. An optimal distillation method should satisfies both criteria.
@@ 4 Experiments
We seek to empirically validate the ability of DSD to guide diffusion-based optimization and generate diverse and high-fidelity outputs. We first analyze our formulation in the simple but informative task of image optimization and then examine text-to-3D generation. We further ablate some key decision decisions and also highlight the applicability for distillation-based single-view reconstruction.
Figure 5: Generation Comparison. We visualize 4 text-to-3D generations from various score distillation methods. We find that DSD is capable of generating high-quality 3D shapes while being more diverse compared to prior methods. Please see supplementary for videos.
@@
We select SDS  [24], HiFA [41] and ASD  [21] to represent the methods that sample random noise, and select Consistent 3D  [36] and SDI  [20] as the approaches that perform score distillation from an ODE perspective. We also include VSD  [35] as it minimizes the KL divergence between the rendering and image distribution, thus promising diverse generation. For fair comparison across methods, we adopt the codebase from SDI (which builds on a commonly adopted implementation [6]) and use a shared one-stage optimization scheme. We only vary the score function across methods (except for VSD, where we use its released code for multi-particle optimization).